OpenStudy (anonymous):

write the quadratic equation in standard form with the given solution set {-4/3,2/5}

OpenStudy (triciaal):

the given solution set means {-4/3,2/5} are the x values (when y = 0) (x+4/3)(x -2/5)= 0

OpenStudy (triciaal):

standard form for the quadratic equation is Ax^2 + Bx+ C = 0 some people use Ax^2 + Bx = C

OpenStudy (triciaal):

multiply everything by the LCD to get rid of fraction

OpenStudy (triciaal):

what did y0u get?